Circular No. 7604

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 F1 (NEAT) E. F. Helin, S. Pravdo, and K. Lawrence, Jet Propulsion Laboratory,

report the discovery of a comet with a faint tail ∼ 40′′ long toward the west-northwest on CCD images taken with the NEAT 1.2-m reflector at Haleakala (discovery observation given below). Additional observations, together with orbital elements (T = 2001 Jan. 21, q = 4.3 AU, i = 19^{o}, P = 15.4 yr) by B. G. Marsden, are given on MPEC 2001-F51. The object appears diffuse on Mar. 28.5 UT CCD images taken by G. J. Garradd, Loomberah, N.S.W. (0.45-m f/5.4 reflector). CCD observations by M. Tichý and M. Kočer at Kleť (0.57-m f/5.2 reflector) on Mar. 29.0 show a diffuse 10′′ coma. P. G. Comba, Prescott, AZ, reports that CCD images taken with a 0.46-m f/4.5 reflector on Mar. 29.3 show a tail in p.a. 285^{o}.

2001 UT  α_{2000}  δ_{2000}  m_{1}
Mar. 24.41577  13^{h}25^{m}41.^{s}76  +17^{o}33′44′′.0  19.7

SUPERNOVA 2001ah IN UGC 6211 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001ah (cf. IAUC 7603), obtained by P. Berlind on Mar. 28.21 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a peculiar type-Ia supernova well before maximum light. The Si II 635.5-nm feature is extremely weak or absent, but the Fe III lines at 430 and 500 nm are present, just as in SN 1991T. There is, however, strong Ca II H and K absorption, implying that SN 2001ah is more like SN 1998ab (cf. IAUC 6858) or SN 1998es (cf. IAUC 7054). Narrow Hα emission from a nearby H II region indicates a recession velocity of 16 800 km/s for the host galaxy. The expansion velocity derived from the minimum of the Ca II line (rest 395.1 nm) is ∼ 20 400 km/s.

SUPERNOVA 2001U IN NGC 5442 A. V. Filippenko, University of California, Berkeley; and A. J. Barth,

Harvard-Smithsonian Center for Astrophysics, report that inspection of a CCD spectrum (range 430–680 nm), obtained on Mar. 29 UT with the Keck-I telescope, reveals that SN 2001U (IAUC 7583) is of type Ia, roughly 2 months past maximum brightness.

2001 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green